OpenStudy (anonymous):

[6.03] What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 4 −2 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

3x + 2y = 8 -----(1) 2x + 3y = 2 -----(2) Multiply equation (1) by 2 and equation (2) by 3 6x+4y=16 6x+9y=6 subtract -5y = 10 y=-2 plug this y into equation (1) 3x+2(-2) = 8 3x-4=8 3x=12 x=4

OpenStudy (anonymous):

thanks i worked it out i got the same answer.